OpenStudy (anonymous):

What is the distance between the points (-5, 2) and (4, 9)? A. square root of 32 B. square root of 82 C. square root of 130 D. square root of 170

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

from -5 to 4 is 9 units right from 2 to 9 is 7 units up by pythagoras your answer is \[\sqrt{9^2+7^2}\]

OpenStudy (misty1212):

if you are a slave to the formula you can use \[d=\sqrt{(x_2-x_2)^2+(y_2-y_1)^2}\] you will get the same thing

OpenStudy (anonymous):

so 81 +289=370 so the square root of 370?

OpenStudy (misty1212):

hmm no

OpenStudy (misty1212):

\(7^2=49\)

OpenStudy (anonymous):

ugh I hate calculators. ok so 81+49?

OpenStudy (misty1212):

yeah easier to do this one in your head yeah, \(\sqrt{130}\)

OpenStudy (misty1212):

or just go with C since it is C at least 87% of the time

OpenStudy (misty1212):

btw your picture is cool who is it?

OpenStudy (anonymous):

one of my drawings thank ya :)

OpenStudy (misty1212):

you drew it! wow!

OpenStudy (anonymous):

yeah I love drawing.

OpenStudy (misty1212):

someday make me one and i will change my picture

OpenStudy (misty1212):

can you make it look like me?

OpenStudy (misty1212):

always wanted cat eyes

OpenStudy (anonymous):

lol sure (:

OpenStudy (misty1212):

i will send you my email

OpenStudy (misty1212):

there sent if you make me one i would love it no hurry, but i could use a new profile picture

OpenStudy (anonymous):

okee dokee. ill draw it as soon as I can. I can send it to you maybe as soon as tonight or maybe a little ealier depending on when I can get done with these problems

OpenStudy (misty1212):

no hurry i am grateful any time